OpenStudy (anonymous):

find the maximum y-value on the graph of y=f(x) f(x)= -x^2 + 2x +1 the maximum y-value is y =

OpenStudy (anonymous):

hi, to find the maximum value you need to differentiate the equation twice, if the 2nd derivative is <0, then you have a maximum point.

OpenStudy (anonymous):

\[\frac{ dy }{dx }= -2x+2\]

OpenStudy (anonymous):

\[\frac{ d^2y }{ dx^2 }=-2\]

OpenStudy (anonymous):

so you can see that your 2nd derivative is a negative value, hence you have a maximum point

OpenStudy (anonymous):

so the maximum value would be -2?

OpenStudy (anonymous):

i think what the question is asking you is to draw the graph and figure the maximum value from the graph

OpenStudy (anonymous):

i am drawing it now i will send you the graph

OpenStudy (anonymous):

there is no graph, it is just asking for a maximum value from that equation

OpenStudy (anonymous):

sorry i was on the phone

OpenStudy (anonymous):

yes i read your question again

OpenStudy (anonymous):

ind the maximum y-value on the graph of y=f(x) f(x)= -x^2 + 2x +1

OpenStudy (anonymous):

your teacher has specified the method

OpenStudy (anonymous):

they want you to figure the answer from the graph

OpenStudy (anonymous):

so this means you have to factorize your equation and sketch it

OpenStudy (anonymous):

so you need to find the vertex of the equation

OpenStudy (anonymous):

i will show you how to do that

OpenStudy (anonymous):

\[y= \frac{ 4ac-b^2 }{ 4 }\]

OpenStudy (anonymous):

sorry i wrote your equation the wrong way

OpenStudy (anonymous):

\[f(x) = -x^{2}+2x+1\]

OpenStudy (anonymous):

can you confirm

OpenStudy (anonymous):

(x+1)(x+1)?

OpenStudy (anonymous):

i know the answer

OpenStudy (anonymous):

but i want you to see how to get that answer

OpenStudy (anonymous):

your y value is -2 but if i ask you to prove that graphically can you do so?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well what you do is that you have to draw it

OpenStudy (anonymous):

as you did you factorize the equation then find the x values

OpenStudy (anonymous):

then you substitute or plug the x values in the main equation to find y

OpenStudy (anonymous):

you can have a look at the graph

OpenStudy (anonymous):

you can draw that in ms word

OpenStudy (anonymous):

i can't come up with the answer, thank you anyways

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the answer is -2 for max y value

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to draw a function in general?

OpenStudy (anonymous):

like making a table of x/y values

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now if you do that you'll see the max y value graphically

OpenStudy (anonymous):

another way to that is the following method

OpenStudy (anonymous):

set x= 0 to see were your graph intersect with the y-axis

OpenStudy (anonymous):

then set y=0 to see where your graph intersect with the x-axis

OpenStudy (anonymous):

to find the maximum value of y, use the vertex equation

OpenStudy (anonymous):

\[y= \frac{ 4ac-b^2 }{ 4 }\]

OpenStudy (anonymous):

and to find x of the maximum y

OpenStudy (anonymous):

\[x=-\frac{ b }{2a }\]

OpenStudy (anonymous):

okay thank you